# Scaling Behaviors of LLM Reinforcement Learning Post-Training: An Empirical Study in Mathematical Reasoning

### v2

@@ 1 Introduction
Based on our analysis across these regimes, we propose a predictive formulation that characterizes the relationship between test loss $L$, model size NN, and resource budget $X$ (where $X$ denotes either Compute $C$ or Data $D$). We find that the scaling behavior can be effectively modeled by a power law that exhibits a log-linear relationship between test loss and resource budget:
@@ 3.1 Compute-Optimal Scaling
To characterize the scaling behavior under computational limits, we first formalize the Compute-Constrained Scenario. Given a fixed computational budget $C$, we seek to identify the optimal model size NN (and the corresponding data allocation $D$) that minimizes the final test loss. This can be expressed as the following constrained optimization problem:
@@
We further analyze learning efficiency term $k_{C}(N)$ in Eq. 6. As Figure 4 shows, $k_{C}(N)$ grows with model size NN, meaning larger models consistently have higher learning efficiency. However, the efficiency gain from model scale is not uniformly linear. Beyond 32B, the increase in $k_{C}(N)$ diminishes, leading to efficiency saturation.
@@ 3.2 Data-Optimal Scaling
In many practical applications, the bottleneck lies not in compute but in the availability of high-quality reasoning data. We define the Data-Constrained Scenario as determining the model size NN that yields the lowest test loss given a limited amount of unique training data $D$:
@@
In the data-constrained setting, we train models with varying parameter counts NN on fixed amounts of unique samples $D$. As shown in Figure 6, larger models consistently achieve lower test loss and higher sample efficiency across both base and instruct variants. The loss–data relationship also follows a log-linear trend and can be modeled by a similar formula addressed in compute scenario:
@@ 3.5 Domain Transfer
Foundational Scaling Laws of Neural Language Models. Foundational scaling studies show language modeling loss follows smooth power-laws in model size NN, data $D$, and compute $C$ (kaplan2020scaling), with compute-optimal training prescribing near lockstep growth of parameters and tokens under fixed FLOPs (hoffmann2022training). Later analyses attribute earlier discrepancies to embedding/non-embedding parameter accounting, last-layer costs, optimizer warmup, and scale-sensitive hyperparameters (pearce2024reconciling; porian2024resolving), while data-centric refinements examine pruning efficiency (sorscher2022beyond), repetition effects (hernandez2022scaling), gzip-based complexity predictors (pandey2024gzip), constrained or synthetic regimes (muennighoff2023scaling; qin2025scaling), and task transfer (e.g., translation) (isik2024scaling). Test-time compute amplification supplies an inference analogue to classical training laws (snell2024scaling).
@@ C.1 FLOPs Calculation Methodology
The computational cost for a LLM is primarily determined by the number of non-embedding parameters (NN) and the number of processed tokens ($T$). The costs for the fundamental operations are:
@@ D.1 Loss Decomposition Model
  • $G(N)$ denotes the model-limited loss, capturing the asymptotic loss floor imposed by finite model capacity NN in the limit of infinite data. It corresponds to the capacity-dependent performance frontier of models with size NN.
  • $\lambda(N)$ denotes the learnable capacity, defined as the maximum achievable reduction in loss that a model of size NN can attain through post-training, beyond its model-limited loss. In RL post-training settings, this term can depend on the pretraining regime, as well as the degree of mismatch between the pretraining and post-training task settings. Conceptually, the learnable capacity is governed by two opposing effects: larger models generally have greater capacity to extract and acquire new knowledge, while simultaneously having already absorbed more information during pretraining, leaving less headroom for additional improvement via RL. As a result, the monotonic dependence of $\lambda(N)$ on model size is non-trivial, and its precise modeling likely requires additional assumptions and empirical characterization.
@@
where $D_{0}(N)$ denotes the characteristic dataset scale at which half of the learnable capacity is realized, and is treated as a NN-dependent parameter.
@@ D.3 Discussion: Effective log–log slope
The resulting $k_{\max}(N)$ depends only on the model size NN through the parameters of the loss decomposition, and is uniformly bounded by $K_{\max}=\beta/2$.
From this perspective, the slope function $k(N)$ adopted in the main analysis (Equation equation 2) can be interpreted as a parsimonious, low-parameter approximation to the effective maximal slope $k_{\max}(N)$ in Equation equation 20. This establishes structural consistency between the two descriptions: the loss decomposition model potentially captures the finer-grained $(N,D)$-dependent behavior, while the main-text formulation summarizes its dominant NN-dependent trend in a compact form.

### v3

@@ 1 Introduction
Based on our analysis across these regimes, we propose a predictive formulation that characterizes the relationship between test loss $L$, model size $N$, and resource budget $X$ (where $X$ denotes either Compute $C$ or Data $D$). We find that the scaling behavior can be effectively modeled by a power law that exhibits a log-linear relationship between test loss and resource budget:
@@ 3.1 Compute-Optimal Scaling
To characterize the scaling behavior under computational limits, we first formalize the Compute-Constrained Scenario. Given a fixed computational budget $C$, we seek to identify the optimal model size $N$ (and the corresponding data allocation $D$) that minimizes the final test loss. This can be expressed as the following constrained optimization problem:
@@
We further analyze learning efficiency term $k_{C}(N)$ in Eq. 6. As Figure 4 shows, $k_{C}(N)$ grows with model size $N$, meaning larger models consistently have higher learning efficiency. However, the efficiency gain from model scale is not uniformly linear. Beyond 32B, the increase in $k_{C}(N)$ diminishes, leading to efficiency saturation.
@@ 3.2 Data-Optimal Scaling
In many practical applications, the bottleneck lies not in compute but in the availability of high-quality reasoning data. We define the Data-Constrained Scenario as determining the model size $N$ that yields the lowest test loss given a limited amount of unique training data $D$:
@@
In the data-constrained setting, we train models with varying parameter counts $N$ on fixed amounts of unique samples $D$. As shown in Figure 6, larger models consistently achieve lower test loss and higher sample efficiency across both base and instruct variants. The loss–data relationship also follows a log-linear trend and can be modeled by a similar formula addressed in compute scenario:
@@ 3.5 Domain Transfer
Foundational Scaling Laws of Neural Language Models. Foundational scaling studies show language modeling loss follows smooth power-laws in model size $N$, data $D$, and compute $C$ (kaplan2020scaling), with compute-optimal training prescribing near lockstep growth of parameters and tokens under fixed FLOPs (hoffmann2022training). Later analyses attribute earlier discrepancies to embedding/non-embedding parameter accounting, last-layer costs, optimizer warmup, and scale-sensitive hyperparameters (pearce2024reconciling; porian2024resolving), while data-centric refinements examine pruning efficiency (sorscher2022beyond), repetition effects (hernandez2022scaling), gzip-based complexity predictors (pandey2024gzip), constrained or synthetic regimes (muennighoff2023scaling; qin2025scaling), and task transfer (e.g., translation) (isik2024scaling). Test-time compute amplification supplies an inference analogue to classical training laws (snell2024scaling).
@@ C.1 FLOPs Calculation Methodology
The computational cost for a LLM is primarily determined by the number of non-embedding parameters ($N$) and the number of processed tokens ($T$). The costs for the fundamental operations are:
@@ D.1 Loss Decomposition Model
  • $G(N)$ denotes the model-limited loss, capturing the asymptotic loss floor imposed by finite model capacity $N$ in the limit of infinite data. It corresponds to the capacity-dependent performance frontier of models with size $N$.
  • $\lambda(N)$ denotes the learnable capacity, defined as the maximum achievable reduction in loss that a model of size $N$ can attain through post-training, beyond its model-limited loss. In RL post-training settings, this term can depend on the pretraining regime, as well as the degree of mismatch between the pretraining and post-training task settings. Conceptually, the learnable capacity is governed by two opposing effects: larger models generally have greater capacity to extract and acquire new knowledge, while simultaneously having already absorbed more information during pretraining, leaving less headroom for additional improvement via RL. As a result, the monotonic dependence of $\lambda(N)$ on model size is non-trivial, and its precise modeling likely requires additional assumptions and empirical characterization.
@@
where $D_{0}(N)$ denotes the characteristic dataset scale at which half of the learnable capacity is realized, and is treated as a $N$-dependent parameter.
@@ D.3 Discussion: Effective log–log slope
The resulting $k_{\max}(N)$ depends only on the model size $N$ through the parameters of the loss decomposition, and is uniformly bounded by $K_{\max}=\beta/2$.
From this perspective, the slope function $k(N)$ adopted in the main analysis (Equation equation 2) can be interpreted as a parsimonious, low-parameter approximation to the effective maximal slope $k_{\max}(N)$ in Equation equation 20. This establishes structural consistency between the two descriptions: the loss decomposition model potentially captures the finer-grained $(N,D)$-dependent behavior, while the main-text formulation summarizes its dominant $N$-dependent trend in a compact form.